1. Give a unit circle argument to support sin(a + 1) =-sin(a) 2. Use the unit circle to solve sin(x) >0 for x = (0,27)Please help explain these two!
Answer #1

Sin(a+1) = -sinld) Sin(A+B)= sinca) COSCB) + COSCA) sin(B) / ( hs Sin(2+1) = sin(a) cos(11) + Cosla) .SinCH) Sinct)=0 COS(T):